Counties, Divisions, or Parishes for 50 States The number of counties, divisions, or parishes for each of the 50 states is given below. Use the data to construct a grouped frequency distribution with 6 classes, a histogram, a frequency polygon, and an ogive. Analyze the distribution. 67 27 15 75 58 64 8 67 159 5 102 44 92 99 105 120 64 16 23 14 83 87 82 114 56 93 16 10 21 33 62 100 53 88 77 36 67 5 46 66 95 254 29 14 95 39 55 72 23 3 Source: World Almanac and Book of Facts.
Answer #1

Minimum : 3, Maximum: 254

Number of classes required: 6

Class width = (maximum -minimum) /6= (254-3)/6=41.83

Hence, the required class width is 42.

Following table shows the grouped data, cumulative frequency :

Class limits LowerUppM Midpoint-(lower+upper)/2 Frequency 45 87 129 Cumulative frequency 19 17 12 24 19 36 48 49 49 50 45 87 129 108 150 192 234 213 255 0 213

Following is the histogram:

Following is the frequency polygon:

Frequency Polygon 20 1 16 14 12 10 6 4 0 0 24 108 150 192 234 276 Classes mid points

Ogive:

Ogive 60 50 40 a 30 5 20 10 0 45 87 129 213 255 Upper limit